Agenda/Minutes of the VIS Meeting on 2020/6/19
2020/6/19 14:00 - 15:00
Participants: Takahashi, Hirata, Fabian, Miyo, Shoda, Sato, Lucia, Tanaka, Washimi, Miyoki, Ishizaki
[By Akutsu] Just for announcement: the 1st draft of the OSEM technical paper has been circulated to some of you; uploaded here; the main topic is evaluations of the sensor and actuator of our OSEM with a central hole widened after the iKAGRA run in 2016, and so the co-authors were selected accordingly by me so far (I'm still increasing the numbers of the names, referring to the past exchanged mails discussing who took which data). One of my concern during the writing is the thermal noise due to eddy current loss in the coil body. The sensor noise of a spare OSEM stocked at Mitaka was recently measured by me, and I got 0.5 nm/rtHz at 1 Hz, and 0.1 nm/rtHz at 10 Hz in the end (evaluated with single-end signaling, though). It would be nice if you can check this is consistent with the output from the installed ones (but I think such evaluations would be out of the scope of this paper). The noise level seems to be dominated by the intensity noise of the LED, so it can be improved with several methods if you want (or, if you would like to improve it more from the viewpoint of vibration-isolation performance).
- For SR baffle installation:
SR2 decay times: following the issue of the mode IP-L 406 mHz, whose damping time was 35 seconds longer with the control system on (klog entry 14510)
I checked the UGF to be 85 mHz (klog entry 14550). Being lower than 406 mHz it should not have an affect according to certain expectations. The OLTF gain at such frequency was 0.22. Still, I put a notch (f=406 mHz, Q=10 and depth 40 dB) in the filter and repeated the measurement. The decay time with the control on decreased by 40 seconds. It became 5 seconds shorter than with the control off. See klog entry 14574.
- The mystery then becomes the mode IP-L at 671 mHz, which is above the UGF and yet the decay time with the control on is 81 seconds shorter than with the control off. In principle the control system shouldn't be doing anything. I will add an notch at 671 mHz just to check the effect.
- I gave feedback to Takahashi-san about tasks towards O4.
- The Mexican Physical Society is holding it's annual conference online because of Covid-19. I submitted an abstract in Spanish. I should check what's CPC's procedure for this.
- We completed the magnetic Damper section of the Tyep-B paper.
Terrence created a library called Kontrol: JGW-T2011764.
- Type-A simulation for cryo-pay upgrade
- Simulnk model update is done.
Need update on actuator efficiency according to the klog
- Miyo-kun summarize some of the noise sources.
- They should be integrated together.
- Type-A modeling for paper writing
- Simulated the spring constant of HLs using COMSOL
- It reproduced the effect of the stranded wires that Yamada-kun showed in his thesis.
- However, obtained spring constant with a realistic parameter is far smaller than the one we need to reproduce the obtained transfer functions.
- Making the circuit of filters for the LVDT combiner.
- Adjustment of inductance took much time.
- Start experiments in the next week.
- Post Obs. tasks are on going. Real-time model was modified in this week. Sensor correction for SRs and SR suspension update are planed in the next week.
- The extended chief meeting was held on 15th. DRFPMI commissioing was scheduled in Jul.-Aug. The tasks towards O4 will be started from Sep.
Proposal on suspension commissioning was shown in the extended chief meeting.
- "Remote/onsite support on VIS troubles by VIS group members with shift style" is requested for the DRFPMI commissioing.
Schedule towards O4 was revised again.
Correction of concerned topics for the evaluation meeting on 22nd (google spreadsheet).
- On 2020/6/26(Fri)